OpenStudy (anonymous):

You want to invest money for your newborn child so that she will have $50,000 for college on her 18th birthday. Determine how much you should invest if the best annual rate that you can get on a secure investment is 5.2% compounded annually. Round your answer to two decimal places.

OpenStudy (mertsj):

A=P(1+r/n)^nt

OpenStudy (johnweldon1993):

so the rate of compounding is 5.2% or .052 use V = Vi (1 + r)^t t = 18 years V = 50000 r = .052 you want to find Vi

OpenStudy (anonymous):

so that would be $50,000=vi (1+0.052)^18

OpenStudy (johnweldon1993):

right....just solve for Vi

OpenStudy (anonymous):

vi=20076.44?

OpenStudy (johnweldon1993):

that's what I get :)

OpenStudy (anonymous):

let me check really quickly..

OpenStudy (anonymous):

yes it works...thank you :D

OpenStudy (johnweldon1993):

not a problem :)